OpenStudy (anonymous):

Find the soloution of this system of eqations. -4y=-64x 4x-4y=-36 i have hd trouble with these kind of quetions can somebod please help me?(:

OpenStudy (anonymous):

If you have a graphing calculator, just put them in and the intersecting point is the answer.

OpenStudy (anonymous):

64x - 4y = 0 4x - 4y = -36 -64x + 4y = 0 4x - 4y = -36 -60x = -36 x = 36/60 x = 3/5